Please help me solve this equation: Given the functions f(x)=2x/x-3 and g(x)=x-2 find a fomula, in simplest form, for (f o g)(x).
Found 2 solutions by stanbon, Theo:
Answer by stanbon(75887)

Given the functions f(x)=2x/x-3 and g(x)=x-2 find a fomula, in simplest form, for
(f o g)(x) = f[g(x)] = f(x-2) = [2(x-2)]/[x-2-3] = (2x-4)/(x-5)
-----
Cheers,
Stan H.
======

Answer by Theo(13342)

f(x) = 2x/(x-3)
g(x) = (x-2)
fog(x) = f(g(x)) = f(x-2) = 2(x-2) / ((x-2)-3)) = (2x-4) / (x-5)
-----
you are replacing x in f(x) with g(x).
since g(x) is equal to (x-2), you are replacing x in f(x) with (x-2).